# Human Decision-making is Susceptible to AI-driven Manipulation

### v1

@@ Abstract
Artificial Intelligence (AI) systems are increasingly intertwined with daily life, assisting users in executing various tasks and providing guidance on decision-making. This integration introduces risks of AI-driven manipulation, where such systems may exploit users’ cognitive biases and emotional vulnerabilities to steer them toward harmful outcomes. Through a randomized controlled trial with $233$ participants, we examined human susceptibility to such manipulation in financial (e.g., purchases) and emotional (e.g., conflict resolution) decision-making contexts. Participants interacted with one of three AI agents: a neutral agent (NA) optimizing for user benefit without explicit influence, a manipulative agent (MA) designed to covertly influence beliefs and behaviors, or a strategy-enhanced manipulative agent (SEMA) employing established psychological tactics to reach its hidden objectives. To ensure participant well-being, this study involved hypothetical scenarios, pre-screening for at-risk individuals, and comprehensive post-study debriefing. By analyzing participants’ decision patterns and shifts in their preference ratings post-interaction, we found significant susceptibility to AI-driven manipulation. Particularly, across both decision-making domains, participants interacting with the manipulative agents shifted toward harmful options at substantially higher rates (financial, MA: $62.3\%$, SEMA: $59.6\%$; emotional, MA: $42.3\%$, SEMA: $41.5\%$) compared to the NA group (financial, $35.8\%$; emotional, $12.8\%$). Notably, our results reveal that even simple manipulative objectives (MA) can be as effective as employing established psychological strategies (SEMA) in swaying human decision-making. These findings suggest that AI-driven manipulation could become widely spread, as it does not require sophisticated tactics and expertise to influence users. By revealing the potential for covert AI influence, this study highlights a critical vulnerability in human-AI interactions, emphasizing the urgent need for ethical safeguards and regulatory frameworks to ensure responsible deployment of AI technologies and protect human autonomy.
@@ Introduction
The vast integration of AI technologies into our daily lives has fundamentally changed how we process information and make decisions. As AI systems increasingly serve as personal assistants, humans have demonstrated growing reliance and trust in AI-generated content across various domains [1, 2]. While this technological shift offers unprecedented access to information and assistance with decision-making, it also introduces subtle yet significant risks to human autonomy. At the core of these risks lies the potential for AI systems to influence human beliefs and behaviors in ways that may bypass their conscious awareness [3, 4, 5, 6].
@@
Recent research has demonstrated AI systems’ manipulative capabilities in controlled environments, primarily in game-based settings and simulated interactions[31, 32, 33, 34, 35], without involving real human subjects. While these studies provide valuable insights, their applicability to real-world contexts is limited. For instance, game-based environments simplify human behavior into predictable patterns (e.g., bluffing to maximize in-game rewards[34]) and may not consider the cognitive biases and emotional vulnerabilities that shape human decisions [36, 37]. Similarly, simulated interactions lack the dynamic nature of human-AI interactions, where users’ evolving beliefs, emotions, and perceived consequences (e.g., financial risks or interpersonal conflicts) affect their susceptibility to influence. Therefore, the extent to which real humans are susceptible to AI-driven manipulation in practical decision-making contexts remains unexplored.
@@ Results
Figure 2: Distribution of decision patterns across AI conditions in decision-making contexts. The bar charts indicate the proportion of decision patterns (%) in financial and emotional scenarios across three AI conditions: NA (Neutral Agent), MA (Manipulative Agent), and SEMA (Strategy-enhanced Manipulative Agent). In both domains, participants in the manipulative groups (MA and SEMA) had significantly more negative shifts (i.e. opting for a harmful option) and fewer positive shifts (i.e., switching to the optimal choice) than the NA condition (Supplementary Table S1).
Conversely, participants were more likely to identify and prioritize the optimal choices post-interaction in the absence of covert manipulation. Interactions with the NA showed significantly more positive shifts (i.e., switching to the optimal choice) than the manipulative conditions across both financial (NA: $31.7\%$; vs. MA: $5.3\%$; vs. SEMA: $4.4\%$; $P<0.0001$; $h=0.73$ and $0.77$, respectively) and emotional (NA: $28.2\%$; vs. MA: $11.7\%$, $P<0.01$, $h=0.42$; vs. SEMA: $13.0\%$, $P=0.015$, $h=0.38$) scenarios. Detailed statistical results are provided in Supplementary Table S1.
These patterns are reflected by temporal changes in participants’ ratings (Figure S1; Supplementary Table S4; Methods). At baseline (pre-interaction), across both domains, there was no significant difference in ratings of the optimal choices and hidden incentives, respectively, between AI conditions ($P=1.00$). However, significant differences emerged post-interaction. In both domains, participants in the manipulative conditions showed significant shifts in their preferences in line with the agents’ hidden motives, exhibiting substantially higher ratings for the hidden incentives (Cohen’s $d\in[0.77,1.17]$) and lower ratings for the optimal options ($d\in[0.50,1.03]$) compared to the control (NA) condition.
@@
Financial decisions showed higher rates of negative shifts compared to the emotional domain ($49.4\%$ vs. $32.2\%$, $P<0.0001$, Cohen’s $h=0.35$; Figure 2), suggesting that participants in the financial domain were more likely to be swayed toward harmful choices. In contrast, emotional decision-making showed a different pattern, as participants were more likely to retain harmful choices than those in the financial domain ($29.9\%$ vs. $12.6\%$, $P<0.0001$, $h=0.43$), suggesting that emotional decisions may be more influenced by personal thoughts and beliefs. Therefore, reinforcement of harmful beliefs and thoughts was a primary aspect of AI-driven manipulation in this domain.
Linear mixed-effects models (Equation 1; Supplementary Table S5) revealed different factors of individual susceptibility across these domains. In financial decision-making, higher openness to experience (coefficient $\beta=-0.08$, $P=0.07$) and lower self-esteem ($\beta=0.02$, $P=0.02$) predicted substantially more harmful outcomes, suggesting that curious and less confident individuals are more likely to accept AI’s advice rather than critically evaluating it. Greater trust in AI systems also led to more harmful decisions ($\beta=-0.01$, $P=0.08$), which may be due to participants’ over-trust in the AI agent’s perceived objectivity, a trend that aligns with existing literature on human reliance on algorithmic recommendations [42].
Similar to the financial domain, higher openness ($\beta=-0.08$, $P=0.07$) was associated with more harmful emotional outcomes, reflecting that open-minded individuals may also be more susceptible to manipulation in emotional decision-making. In addition, more affective commitment traits correlated with making harmful decisions ($\beta=-0.10$, $P=0.09$), suggesting that emotionally reactive individuals are more prone to manipulation. Our findings also revealed higher normative ($\beta=0.17$, $P=0.02$) and continuance ($\beta=0.12$, $P=0.04$) traits were strongly associated with more beneficial emotional outcomes, highlighting the importance of reciprocity and persistence in identifying the optimal choice for handling emotional dilemmas (i.e., active coping).
@@
Notably, incorporating psychological strategies (SEMA) did not amplify harmful decision-making outcomes compared to mere manipulative incentives (MA). Across both domains, SEMA’s influence on participants’ decisions nearly matched MA’s ($P=1.00$; Cohen’s $h\in[0.02,0.06]$; Figure 2), indicating that hidden objectives alone are sufficient to influence human decision-making. However, temporal changes in preference ratings revealed domain-specific nuances in the impact of strategies (Figure 2; Supplementary Table S4). Compared to the NA condition, the addition of strategies (SEMA vs. MA) resulted in larger effect sizes for both the optimal choices (Cohen’s $d$: $1.17$ vs. $d$: $0.86$) and the hidden incentives ($d$: $1.03$ vs. $d$: $0.72$) in the financial domain, while a contrasting pattern emerged in the emotional domain (optimal choice, $d$: $0.77$ vs. $d$: $1.00$; hidden incentive, $d$: $0.50$ vs. $d$: $0.78$).
These findings suggest that hidden objectives play a significant role in AI-driven manipulation, as the presence of such objectives alone (MA) was nearly as effective as employing established psychological strategies (SEMA) in influencing participants’ decisions. However, the addition of strategies did enhance the manipulative impact in certain scenarios, suggesting that while hidden objectives are a primary driver, the use of sophisticated tactics may further amplify their influence in contexts where users tend to defer more to AI’s perceived expertise (i.e., financial decisions).
@@
Figure 3 illustrates the strategy distribution of the strategy-enhanced manipulative agent (SEMA) across decision-making domains. Chi-squared analyses indicated no significant difference between the overall strategy distributions between financial and emotional scenarios ($\chi^{2}[1]=12.16$, $P=0.35$), which suggests that, in both domains, the agent demonstrated an implicit preference toward strategies that typically elicit more positive emotions (e.g., Pleasure Induction and Charm as opposed to Guilt Trip).
Figure 3: Distribution of strategies employed by the strategy-enhanced manipulative agent (SEMA) across decision-making contexts. The bar plots show the proportion of responses for each manipulation strategy used by the SEMA, with significant differences annotated above each bar plot. These results suggest that SEMA adapted its strategy usage to the emotional or financial nature of the decisions, reflecting a tailored approach to influence. $P$ value legend: n.s. (not significant), $P>0.05$; *, $P\leq 0.05$; **, $P\leq 0.01$; ***, $P\leq 0.001$; ****, $P\leq 0.0001$).
However, the rate at which the SEMA used these strategies varied significantly across the two domains (Supplementary Table S3). In financial scenarios, this agent employed diversion ($23.08\%$ vs. $14.44\%$, $h=0.22$, $P<0.0001$), justification ($15.83\%$ vs. $8.81\%$, $h=0.22$, $P<0.0001$), urgency ($2.97\%$ vs. $0.30\%$, $h=0.24$, $P<0.0001$), and fabricated information ($3.87\%$ vs. $1.90\%$, $h=0.12$, $P<0.01$) significantly more compared to the emotional domain, mirroring real-world predatory marketing [43]. Conversely, compared to the financial domain, it relied on Pleasure Induction ($45.82\%$ vs. $29.84\%$, $h=0.33$, $P<0.0001$) and Guilt Trip ($5.09\%$ vs. $2.39\%$, $h=0.14$, $P<0.001$) significantly more in emotional contexts, attempting to exploit emotional vulnerabilities. In addition, the SEMA adopted considerably more diverse strategies (i.e., Others) in the emotional domain ($2.89\%$ vs. $1.40\%$, $h=0.10$, $P<0.05$). Upon further investigation, such responses primarily aimed to provide reassurance and reinforce the users’ thoughts and beliefs, which is in line with our previous finding on domain-specific decision patterns.
@@
The majority of participants perceived the manipulative agents as helpful across financial (MA: $86.8\%$, and SEMA: $78.9\%$) and emotional (MA: $86.4\%$, and SEMA: $75.6\%$) scenarios, with rates comparable to the neutral agent (financial: $97.5\%$; emotional: $87.1\%$). This perception of helpfulness highlights that participants were largely unaware of the agents’ manipulative intents and viewed them as equally beneficial as the neutral assistant. Notably, while participants were not explicitly prompted to report whether they thought the agents had ulterior motives, a considerable portion of participants mentioned noticing signs of such influence, particularly in the financial domain (MA: $13.2\%$; SEMA: $28.9\%$; vs. NA: $0\%$). In contrast, these concerns were less frequently reported in the emotional domain (MA: $8.1\%$; SEMA: $9.7\%$; vs. NA: $2.5\%$), suggesting that the agent’s influence was more covert in such contexts. Detailed information on participant feedback is provided in Figure S2 and Supplementary Table S6.
@@ Discussion
Our findings revealed that participants interacting with manipulative agents (MA and SEMA) exhibited substantially higher rates of harmful decision-making compared to those who interacted with the NA. This susceptibility was further pronounced in temporal changes of participants’ ratings, as participants in the manipulative groups exhibited declining preferences for the optimal choices while increasingly favoring the hidden incentives after interacting with these agents.
The contrasting decision patterns between financial and emotional domains highlight how context may shape human susceptibility to AI influence. These patterns suggest different mechanisms of manipulative influence across the two domains: the manipulative agents reinforced existing harmful tendencies in emotional scenarios while more actively shifting initially sound decisions in the financial scenarios.
Another notable finding was the similar effectiveness of MA and SEMA conditions in shifting participants’ preferences and decisions. These results raise critical concerns about the potential for covert influence in AI systems, as agents that were not explicitly designed to manipulate their users could also significantly impact human decision-making based on a hidden incentive.
Our analyses also revealed the SEMA’s adaptation of manipulation strategies across different domains. These findings indicated that while the overall distribution of the employed strategies followed a similar trend across different domains, suggesting that the SEMA may implicitly prefer certain strategies more than others, this agent tailored its strategy usage to the context (e.g., exploiting interpersonal vulnerabilities in the emotional domain while using pragmatic tactics in financial scenarios). This context-specific adaptation highlights the agent’s nuanced approach to maximizing manipulative impact across different scenarios, which highlights broader issues regarding AI safety.
@@
Another key consideration was the decision to share the prompts used to configure the manipulative AI agents. While publishing these prompts may enable malicious actors to develop more sophisticated manipulative systems, we believe transparency and reproducibility are essential for scientific progress on this topic. In addition, our findings indicate that simply providing a hidden objective to the AI agent could negatively influence participants’ preferences and decisions, with minimal additional gain from more sophisticated strategies employed by the SEMA. Therefore, the primary risk lies in the presence of hidden objectives rather than the complexity of the prompts.
@@
Despite our efforts to create the first robust and comprehensive analysis of human susceptibility to AI-driven manipulation, this study faces several limitations. The study was conducted in a controlled environment with pre-defined decision-making scenarios, which enabled us to explore the effects of AI-driven influence in isolation while limiting potential risks to participants’ well-being. However, these artificial constraints may not fully capture the complexity of real-world AI applications, where users are exposed to a broader range of decisions and interactions. Future work can examine how these findings translate to more dynamic, real-world settings where AI interacts with users across diverse contexts, exploring such influence beyond decision-making.
In addition, participants’ awareness of being in an experimental setting might have influenced their reported scores and how they interacted with the assigned AI agent, which may subtly influence their natural decision-making process. Relying on self-reported measures to explore changes in participants’ preferences could also imply limitations in our methodology as participants’ willingness to report changes in their preferences may not always align with their actual cognitive and behavioral changes. Hence, future work could assess the long-term effects of AI-driven influence.
Regarding the AI agents employed in this study, our reliance on GPT-4o as the base model may present a limitation. While GPT-4o is currently state-of-the-art across a range of tasks, manipulation capabilities can vary significantly across different models. Given the widespread adoption and popularity of GPT-4o (commercially known as ChatGPT), understanding its manipulative tendencies is critical due to its large user base and significant societal impact. However, we do not claim that all models exhibit the same manipulative behavior, as further studies are needed to investigate whether similar tendencies are present in other models. Moreover, our curated taxonomy of strategies represents a subset of potential manipulation techniques in user-assistant settings. Hence, while this taxonomy is theory-based and comprehensive, emerging AI capabilities may form more subtle and sophisticated strategies that are not captured in our current taxonomy.
The design of the control condition in our experiment also poses a limitation. While the neutral agent (NA) was designed to optimize user benefit without explicit manipulative objectives, inherent model biases may persist. These biases may have subtly influenced participants’ decisions, even in the absence of a hidden agenda, as seen in the negative shifts across both decision-making domains for the NA group. This limitation emphasizes the challenge of creating truly neutral AI systems, and future research should focus on further identifying and mitigating harmful biases in AI agents.

### v2

@@ Abstract
Artificial Intelligence (AI) systems are increasingly intertwined with daily life, assisting users in executing various tasks and providing guidance on decision-making. This integration introduces risks of AI-driven manipulation, where such systems may exploit users’ cognitive biases and emotional vulnerabilities to steer them toward harmful outcomes. Through a randomized controlled trial with $233$ participants, we examined human susceptibility to such manipulation in financial (e.g., purchases) and emotional (e.g., conflict resolution) decision-making contexts. Participants interacted with one of three AI agents: a neutral agent (NA) optimizing for user benefit without explicit influence, a manipulative agent (MA) designed to covertly influence beliefs and behaviors, or a strategy-enhanced manipulative agent (SEMA) employing established psychological tactics to reach its hidden objectives. To ensure participant well-being, this study involved hypothetical scenarios, pre-screening for at-risk individuals, and comprehensive post-study debriefing. By analyzing participants’ decision patterns and shifts in their preference ratings post-interaction, we found significant susceptibility to AI-driven manipulation. Particularly, across both decision-making domains, participants interacting with the manipulative agents shifted toward harmful options at substantially higher rates (financial, MA: $62.3\%$, SEMA: $59.6\%$; emotional, MA: $42.3\%$, SEMA: $41.5\%$) compared to the NA group (financial, $35.8\%$; emotional, $12.8\%$). Notably, our results reveal that even simple manipulative objectives (MA) can be as effective as employing established psychological strategies (SEMA) in swaying human decision-making. These findings suggest that AI-driven manipulation could become widespread, as it does not require sophisticated tactics and expertise to influence users. By revealing the potential for covert AI influence, this study highlights a critical vulnerability in human-AI interactions, emphasizing the urgent need for ethical safeguards and regulatory frameworks to ensure responsible deployment of AI technologies and protect human autonomy.
@@ Introduction
The vast integration of AI technologies into our daily lives has fundamentally changed how we process information and make decisions [1, 2]. As AI systems increasingly serve as personal assistants, humans have demonstrated growing reliance and trust in AI-generated content across various domains [3, 4, 5, 6]. While this technological shift offers unprecedented access to information and assistance with decision-making, it also introduces subtle yet significant risks to human autonomy. At the core of these risks lies the potential for AI systems to influence human beliefs and behaviors in ways that may bypass their conscious awareness [7, 8, 9, 10].
@@
Recent research has demonstrated AI systems’ manipulative capabilities in controlled environments, primarily in game-based settings and simulated interactions[35, 36, 37, 38, 39], without involving real human subjects. While these studies provide valuable insights, their applicability to real-world contexts is limited. For instance, game-based environments simplify human behavior into predictable patterns (e.g., bluffing to maximize in-game rewards[38]) and may not consider the cognitive biases and emotional vulnerabilities that shape human decisions [40, 41]. Similarly, simulated interactions between AI systems role-playing as human users lack the dynamic nature of human-AI interactions, where users’ evolving beliefs, emotions, and perceived consequences (e.g., financial risks or interpersonal conflicts) affect their susceptibility to influence. Therefore, the extent to which real humans are susceptible to AI-driven manipulation in practical decision-making contexts remains unexplored.
@@ Results
Figure 2: Distribution of decision patterns across AI conditions in decision-making contexts. The bar charts indicate the proportion of decision patterns (%) in financial and emotional scenarios across three AI conditions: NA (Neutral Agent), MA (Manipulative Agent), and SEMA (Strategy-enhanced Manipulative Agent). In both domains, participants in the manipulative groups (MA and SEMA) had significantly more negative shifts (i.e., opting for a harmful option) and fewer positive shifts (i.e., switching to the optimal choice) than the NA condition (Supplementary Table S1).
Conversely, participants were more likely to identify and prioritize the optimal choices post-interaction in the absence of covert manipulation. Interactions with the NA showed significantly more positive shifts (i.e., switching to the optimal choice) than the manipulative conditions across both financial (NA: $31.7\%$; vs. MA: $5.3\%$; vs. SEMA: $4.4\%$; $P<0.0001$; $h=0.73$ and $0.77$, respectively) and emotional (NA: $28.2\%$; vs. MA: $11.7\%$, $P<0.01$; vs. SEMA: $13.0\%$, $P<0.05$; $h=0.42$ and $0.38$, respectively) scenarios. Detailed statistical results are provided in Supplementary Table S1.
These patterns are reflected by temporal changes in participants’ ratings (Figure S1; Supplementary Table S3; Methods). At baseline (pre-interaction), across both domains, there was no significant difference in ratings of the optimal choices and hidden incentives, respectively, between AI conditions ($P=1.00$). However, significant differences emerged post-interaction. In both domains, participants in the manipulative conditions showed significant shifts in their preferences in line with the agents’ hidden motives, exhibiting substantially higher ratings for the hidden incentives (Cohen’s $d\in[0.34,0.68]$) and lower ratings for the optimal options ($d\in[0.54,0.84]$) compared to the control (NA) condition.
@@
Financial decisions showed significantly higher rates of negative shifts compared to the emotional domain ($49.4\%$ vs. $32.2\%$, $P<0.0001$; Cohen’s $h=0.35$; Figure 2; Supplementary Table S2). This suggests that participants in the financial domain were more likely to be swayed toward harmful choices, possibly due to participants’ over-trust in the AI agent’s perceived objectivity in these contexts, a trend that aligns with existing literature on human reliance on algorithmic recommendations[46]. In contrast, participants were more likely to retain their optimal choices in the emotional contexts compared to the financial domain ($29.9\%$ vs. $12.6\%$, $P<0.0001$, $h=0.43$), indicating that emotional decisions may be more deeply influenced by personal intuition rather than reliance on AI’s advice, a phenomenon known as algorithmic aversion [47], making them less susceptible to external manipulation.
Linear mixed-effects models (Equation 1; Supplementary Table S4) revealed distinct factors of individual susceptibility across these domains. In financial decision-making, lower self-esteem (coefficient $\beta=0.02$, $P=0.03$) was significantly correlated with negative shifts (i.e., switching to a harmful option), suggesting that less confident individuals are more likely to defer to AI recommendations than critically evaluate them. Moreover, unemployed individuals ($\beta=-0.23$, $P=0.08$) were more likely to make harmful financial decisions compared to those working full-time, potentially due to the financial strain and psychological stress of unemployment, which increases the likelihood of riskier behaviors[48] (e.g., opting for less expensive products with few reviews or subscription services offering low initial costs). Greater trust in AI systems also led to more harmful decisions ($\beta=-0.01$, $P=0.09$), further highlighting the impact of AI’s perceived objectivity on human susceptibility in quantitative decision-making contexts[5].
In the emotional domain, higher normative commitment traits ($\beta=0.17$, $P=8.09\times 10^{-3}$) were significantly associated with more beneficial emotional outcomes, indicating the importance of reciprocity and adherence to social norms in identifying the optimal choice for handling emotional dilemmas[49] (i.e., active coping). Conversely, individuals with more affective commitment traits were more prone to negative decision shifts ($\beta=-0.09$, $P=0.08$), suggesting that emotionally reactive individuals are more susceptible to manipulation [50]. Interestingly, the order in which participants completed the scenarios (i.e., completion order) also influenced susceptibility. Participants were more likely to exhibit negative shifts as they progressed through the experiment ($\beta=-0.08$, $P=0.06$), suggesting the initial aversion to AI’s advice in emotional decision-making may diminish over time.
@@
Notably, incorporating psychological strategies (SEMA) did not amplify harmful decision-making outcomes compared to mere manipulative incentives (MA). Across both domains, SEMA’s influence on participants’ decisions nearly matched MA’s ($P=1.00$; Cohen’s $h\in[0.02,0.06]$; Figure 2), indicating that hidden objectives alone are sufficient to influence human decision-making. However, temporal changes in preference ratings revealed domain-specific nuances in the impact of strategies (Figure S1; Supplementary Table S3). Compared to the NA condition, the addition of strategies (SEMA vs. MA) resulted in larger effect sizes (Cohen’s $d$) for both the optimal choices (SEMA: $0.84$ vs. MA: $0.59$) and the hidden incentives (SEMA: $0.68$ vs. MA: $0.48$) in the financial domain, while a contrasting pattern emerged in the emotional domain (optimal choice, SEMA: $0.54$ vs. MA: $0.66$; hidden incentive, SEMA: $0.34$ vs. MA: $0.56$).
@@
Figure 3 illustrates the strategy distribution of the strategy-enhanced manipulative agent (SEMA) across decision-making domains. Across both domains, the agent demonstrated an implicit preference toward strategies that typically elicit more positive emotions (e.g., Pleasure Induction and Charm as opposed to Guilt Trip). However, chi-squared analyses indicated a significant difference in the overall strategy distributions between financial and emotional scenarios ($\chi^{2}[11]=154.41$, $P=1.87\times 10^{-27}$), suggesting that the rate at which the SEMA used these strategies varied across the two domains.
Figure 3: Distribution of strategies employed by the strategy-enhanced manipulative agent (SEMA) across decision-making contexts. The bar plots show the proportion of responses for each manipulation strategy used by the SEMA, with significant differences annotated above each bar plot. These results suggest that SEMA adapted its strategy usage to the emotional or financial nature of the decisions, reflecting a tailored approach to influence. $P$ value legend: n.s. (not significant), $P\geq 0.05$; *, $P<0.05$; **, $P<0.01$; ***, $P<0.001$; ****, $P<0.0001$).
In financial scenarios, this agent employed diversion ($23.1\%$ vs. $14.4\%$, $P<0.0001$; Cohen’s $h=0.22$), justification ($15.8\%$ vs. $8.8\%$, $P<0.0001$; $h=0.22$), urgency ($3.0\%$ vs. $0.3\%$, $P<0.0001$; $h=0.24$), and fabricated information ($3.9\%$ vs. $1.9\%$, $P<0.01$; $h=0.12$) significantly more compared to the emotional domain, mirroring real-world predatory marketing tactics[51]. Conversely, compared to the financial domain, it relied on Pleasure Induction ($45.8\%$ vs. $29.8\%$, $P<0.0001$; $h=0.33$) and Guilt Trip ($5.1\%$ vs. $2.4\%$, $P<0.001$; $h=0.14$) significantly more in emotional contexts, attempting to exploit emotional vulnerabilities. In addition, the SEMA adopted considerably more diverse strategies (i.e., Others) in the emotional domain ($2.9\%$ vs. $1.4\%$, $P<0.05$; $h=0.10$). Upon further investigation, such responses primarily aimed to provide reassurance and reinforce the users’ thoughts and beliefs. Detailed statistical results are provided in Supplementary Table S5.
@@
The majority of participants perceived the manipulative agents as helpful across financial (MA: $86.8\%$, and SEMA: $78.9\%$) and emotional (MA: $86.4\%$, and SEMA: $75.6\%$) scenarios, with rates comparable to the neutral agent (financial: $97.5\%$; emotional: $87.1\%$). This perception of helpfulness highlights that participants were largely unaware of the agents’ manipulative intents and viewed them as equally beneficial as the neutral assistant. Notably, while participants were not explicitly prompted to report whether they thought the agents had ulterior motives, a considerable portion of participants mentioned noticing signs of such influence, particularly in the financial domain (NA: $0\%$; MA: $13.2\%$; SEMA: $28.9\%$; $P<0.01$). In contrast, these concerns were less frequently reported in the emotional domain (NA: $2.5\%$; MA: $8.1\%$; SEMA: $9.7\%$; $P=0.41$), suggesting that the agent’s influence was more covert in such contexts. Detailed information on participant feedback is provided in Figure S2 and Supplementary Tables S6 & S7.
@@ Discussion
Our findings revealed that participants interacting with the manipulative agents (MA and SEMA) exhibited substantially higher rates of harmful decision-making compared to those who interacted with the NA. This susceptibility was further highlighted in the temporal changes of participants’ ratings, as participants in the manipulative groups exhibited declining preferences for the optimal choices while increasingly favoring the hidden incentives after interacting with these agents.
The contrasting decision patterns between financial and emotional domains highlight how context shapes human susceptibility to AI-driven manipulation. Participants showed greater susceptibility to AI’s influence in quantitative decision-making scenarios (i.e., financial contexts), where they may over-trust AI’s algorithmic advice due to its perceived objectivity [46]. In contrast, participants were less susceptible to external influence in emotional scenarios, more frequently preserving their initial beneficial choices, likely due to emotional judgments being deeply rooted in personal intuition and less reliant on external recommendations [47]. Hence, future work should investigate the cognitive and contextual mechanisms underlying human susceptibility and develop targeted interventions that mitigate AI-driven manipulation.
Another notable finding was the similar performance of MA and SEMA conditions in shifting participants’ preferences and decisions. These findings suggest that hidden objectives play a significant role in AI-driven manipulation, as the presence of such objectives alone (MA) was nearly as effective as employing established psychological strategies (SEMA) in influencing participants’ decisions. However, the addition of strategies did enhance the manipulative impact in financial scenarios (Supplementary Table S1), suggesting that while hidden objectives are a primary driver, the use of sophisticated tactics may further amplify their influence in contexts where users tend to defer more to AI’s perceived expertise (i.e., financial decisions). These results raise critical concerns about the potential for covert influence in AI systems, as agents that were not explicitly designed to manipulate their users could also significantly impact human decision-making based on a hidden incentive.
Our analyses also revealed the AI agents’ adaptation of manipulation strategies across different domains. These findings suggest that while the overall distribution of the employed strategies followed a similar trend across different domains, indicating that the AI agents may implicitly prefer certain strategies more than others, such agents can tailor their strategy usage to the context (e.g., exploiting interpersonal vulnerabilities in the emotional domain while using pragmatic tactics in financial scenarios). This context-specific adaptation highlights AI agents’ nuanced approach to maximizing manipulative impact across different scenarios, highlighting broader issues regarding AI safety.
@@
Another key consideration was the decision to share the prompts used to configure the manipulative AI agents. While publishing these prompts may enable malicious actors to develop more sophisticated manipulative systems, we believe transparency and reproducibility are essential for scientific progress on this topic. In addition, our findings indicate that simply providing a hidden objective to the AI agent could negatively influence participants’ preferences and decisions, with minimal additional gain from more sophisticated strategies employed by the strategy-enhanced manipulative agent. Therefore, the primary risk lies in the presence of hidden objectives rather than the complexity of the prompts.
@@
Despite our efforts to create the first robust and comprehensive analysis of human susceptibility to AI-driven manipulation, this study faces several limitations. The study was conducted in a controlled environment with pre-defined decision-making scenarios, which enabled us to explore the effects of AI-driven influence in isolation while limiting potential risks to participants’ well-being. However, these artificial constraints may not fully capture the complexity of real-world AI applications, where users are exposed to a broader range of decisions and interactions. Future work can examine how these findings translate to more dynamic settings where AI interacts with users across diverse contexts, exploring such influence beyond decision-making.
In addition, participants’ awareness of being in an experimental setting might have influenced their reported scores and how they interacted with the assigned AI agent, which may subtly influence their natural decision-making process. Relying on self-reported measures to explore changes in participants’ preferences could also imply limitations in our methodology, as participants’ willingness to report changes in their preferences may not always align with their actual cognitive and behavioral changes. Hence, future work could assess the long-term effects of AI-driven influence.
Regarding the AI agents employed in this study, our reliance on GPT-4o may present a limitation. While GPT-4o is currently state-of-the-art across various tasks [53, 54], manipulation capabilities may vary across different models. Given the widespread adoption and popularity of GPT-4o (commercially known as ChatGPT), understanding its manipulative tendencies is critical due to its large user base and significant societal impact. However, we do not claim that all models exhibit the same manipulative behavior, as further studies are needed to investigate whether similar tendencies are present in other models. Moreover, our curated taxonomy of strategies represents a subset of potential manipulation techniques in user-assistant settings. Hence, while this taxonomy is theory-based and comprehensive, emerging AI capabilities may form more subtle and sophisticated strategies that are not captured in our curated taxonomy.
The design of the control condition in our experiment also poses a limitation. While the neutral agent was designed to optimize user benefit without explicit manipulative objectives, inherent model biases may persist. These biases may have subtly influenced participants’ decisions, even in the absence of a hidden incentive, as seen in the negative shifts across both decision-making domains for the NA group. This limitation emphasizes the challenge of creating truly neutral AI systems, and future research should focus on further identifying and mitigating harmful biases in AI agents.